OpenStudy (anonymous):

let f(x)=x/x-4. find a function y=g(x) so that (fog)(x)=x (f round g) any help?

OpenStudy (goformit100):

just diffrentiate it

OpenStudy (anonymous):

how? i am new at this round thing... can u give me an ex?

OpenStudy (ajprincess):

Let \(f(x)=x+1\) and \(g(x)=x^2\). \(fog(x)=f(g(x))\). To find \(fog(x)\) put g(x) in place of x in f(x). That is \(f(g(x))=(g(x))+1\) \(=x^2+1\) Getting it @aboulmich2?

OpenStudy (anonymous):



OpenStudy (goformit100):

yes

OpenStudy (anonymous):

yes this i got it but in reverse if i have only f(x) how can i find g(x)?

OpenStudy (anonymous):

i mean how to solve the question i wrote in details plzz

OpenStudy (ajprincess):

f(x)=x/x-4 and ur g(x)=y. To find f(g(x)) u wil put g(x) in place of x in f(x) right?

OpenStudy (anonymous):

right

OpenStudy (ajprincess):

f(g(x))=y/y-4 Also u r given that f(g(x))=x. So x=y/y-4 Nw solve for y to find g(x)

OpenStudy (anonymous):

okayy i got it now thanks everyone

OpenStudy (ajprincess):

yw:)